OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (-3, -4) and (0, 2)?

OpenStudy (anonymous):

Slope Formula: 2--2/0--3=4/3 Point Slope Form \[y-y _{1}=m(x-x _{1})\] y--4=4/3(x--3) y+4=4/3x+4 y=4/3x

OpenStudy (anonymous):

These are the answers choices tho (y - 8) = 2(x - 3) (y + 3) = -2(x + 4) (y - 2) = -2(x + 4) (y + 4) = 2(x + 3)

OpenStudy (anonymous):

D

OpenStudy (the_fizicx99):

@sapphire47182 please refrain from giving away answers.

OpenStudy (anonymous):

@sapphire47182 thanks for explaining how to solve it and @tHe_FiZiCx99 , I had a feeling it would be D because I looked at the way she explained it, but I wasn't sure

OpenStudy (the_fizicx99):

It's fine, though giving out answers is against the CoC.

OpenStudy (anonymous):

@tHe_FiZiCx99 I understand